What is the value of x when 1.0E-25 =(3x)^3(2x)^2
Answer by Alan3354(69443)

What is the value of x when 1.0E-25 =(3x)^3(2x)^2
1.0E-25 = 27x^3*4x^2 = 108x^5
x^5 = 1000E-28/108 =~ 9.259E-28
x^5 = 925.9E-30
x = 3.92E-6